Variant 5.

A worker gets wages (salary) from a capitalist. What function will money play in this case? What functions of money do you know? (5 points).

 The productivity of labour has increased twice. What will happen with the volume of output, value of one commodity, the value of total output? What is the difference between labour productivity and labour intensity? (5 points).

The price of commodity has increased from 0,5 to 1 dol. and volume of demand has decreased from 10 to 5 units. Determine the coefficient of price elasticity and comment the result. (10 points).


1
Expert's answer
2021-12-06T10:55:28-0500

Task #273663


Money will be used as a medium of exchange for services offered if a worker receives a salary from the capitalist .The function of money will be a medium of exchange in this case.


Other Functions of money

  1. Unit of account - money serves as a standard numerical unit of measurement of market value of goods, services and other transactions.
  2. Money acts as a store of value - To act as a store of value money should be reliably saved, stored, retrieved and be predictably usable as a medium of exchange when retrieved. It's value must remain stable over time.
  3. Money is used as a standard of deferred payments - where it is an acceptable way to settle a a debt.


If the productivity of labour increases twice, the volume of output will increase .


The value of one commodity in terms of price will reduce because with the increased output there will be more supply of the commodity in the market which will result in reduced value of the commodity.


The value of the total output will increase with the increasing productivity of labour because the production input of labour is being used efficiently to produce more goods which will result in increased consumption.


Difference between labour productivity and labour intensity is as follows:


labour productivity measures the total volume of output produced per unit of labour measured in terms of number of employees during a given time reference period while labour intensity measures the relative proportion of labour compared to capital used in any given process.


Calculation of coefficient of price elasticity.


"P1=0.5\\P2=1\\ Q1=10\\ Q2=5"

Using mid point formula.


"Ep=\frac{Q2-Q1}{\frac{Q1+Q2}{2}}÷\frac{P2-P1}{\frac{P1+P2}{2}}"


"Q2-Q1=5-10=-5\\\frac{Q1+Q2}{2}=\frac{10+5}{2}=7.5\\P2-P1=1-0.5=\$0.5\\\frac{P1+P2}{2}=\frac{1+0.5}{2}=0.75"

"Ep=\frac{-5}{7.5} ÷\frac{0.5}{0.7}= -0.93"


Generally economists ignore the negative sign and price elasticity of demand coefficient is provided in absolute values.


"Ep=0.93"


"Ep<1"

Indicating that the good is inelastic